A guitar factory has a cost of production C(x)=75x+50000. If the company needs to break after 150 units sold, at what price should they sell each guitar to break even?
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52832)

.
75x + 50000 = 150*x


50000 = 150x - 75x


50000 = 75x


x = 50000/75 = 666.67.


ANSWER.  666.67, rounded to nearest greatest cent.


Answer by Boreal(15235)

C=R, x is the price per unit.
so 75x+50000=150x
50000=75 x
x=$666.67